# Relativistic framework for high-precision GNSS-based navigation in cislunar space

### v1

@@ II.6.2 Inverse transformations: proper time $\rightarrow$ TDB
III Processing the GNSS data from the BCRS
In this Section, we discuss the validation of the position, velocity, and acceleration transformations outlined previously, now specifically tailored for GNSS spacecraft. Our focus is on evaluating the precision of spacecraft positioning achieved by processing data within the BCRS. This analysis aims to demonstrate the transformation’s effectiveness in enhancing the accuracy of GNSS spacecraft navigation.
Before proceeding, we will present the equations of motion for both coordinate systems and discuss the relevant dynamics. This overview is crucial for understanding how these transformations impact the positioning and navigation.
@@ III.1.1 BCRS metric in the mass monopole approximation
where the indices $j$ and $k$ refer to the $N$ bodies, with $k$ including body $i$, whose motion is being investigated. The gravitational constant for body $j$, denoted as $\mu_{j}$, is defined as $\mu_{j}=Gm_{j}$, where $G$ is the universal Newtonian gravitational constant, and $m_{j}$ is the isolated rest mass of body $j$. Additionally, the vector $\boldsymbol{\mathrm{r}}_{i}$ represents the barycentric radius-vector of this body. The vector $\boldsymbol{\mathrm{r}}_{ij}=\boldsymbol{\mathrm{r}}_{\tt J}-\boldsymbol{\mathrm{r}}_{i}$ points from body $i$ to body $j$, $r_{ij}=|\boldsymbol{\mathrm{r}}_{\tt J}-\boldsymbol{\mathrm{r}}_{i}|$, and the vector $\boldsymbol{\mathrm{n}}_{ij}=\boldsymbol{\mathrm{r}}_{ij}/r_{ij}$ is the unit vector along this direction.
@@ III.1.3 Light time equation
$t_{2}-t_{1}=\frac{r_{12}}{c}+(1+\gamma)\sum_{\tt B}\frac{GM_{\tt B}}{c^{3}}\ln\left[\frac{r_{1}^{\tt B}+r_{2}^{\tt B}+r_{12}^{\tt B}}{r_{1}^{\tt B}+r_{2}^{\tt B}-r_{12}^{\tt B}}\right]+{\cal O}(c^{-5}),$ (64)
where $t_{1}$ refers to the signal transmission time, and $t_{2}$ refers to the signal reception time. Additionally, $r_{1}$ and $r_{2}$ are the barycentric positions of the transmitter and receiver, respectively, while $r_{12}$ represents their spatial separation.
@@
Frequency transfer observable at $\mathcal{O}(c^{-3})$
@@
where the ratio $q_{A}/q_{B}$ follows from differentiating the light-time (our Eq. (64)) with respect to emission time, retaining terms through $\mathcal{O}(c^{-3})$ [36, 37]. For the two-way (uplink+downlink) configuration used in precise time/frequency comparisons, a convenient form is
$\frac{\nu_{\tt B}}{\nu_{\tt A}}=\frac{1}{2}\left(\frac{\nu_{\tt B}}{\nu_{\tt B^{\prime}}}\right)_{\!\text{station}}+\Delta_{\tt AB}+\frac{1}{2}+\mathcal{O}(c^{-4}),$ (68)
with
$\Delta_{\tt AB}$ $=\frac{1}{c^{2}}\!\left(U_{\tt AB}-{\textstyle\frac{1}{2}}v_{\tt AB}^{2}-(\mathbf{R}_{\tt AB}\cdot\mathbf{a}_{\tt B})\right)\!\left(1+\frac{(\mathbf{N}_{\tt AB}\cdot\mathbf{v}_{\tt AB})}{c}\right)-\frac{R_{\tt AB}}{c^{3}}\Big((\mathbf{v}_{\tt A}\cdot\mathbf{a})_{\tt B}-R_{\tt AB}\,b_{\tt B}-2(\mathbf{v}_{\tt B}\cdot\mathbf{a}_{\tt B})+(\mathbf{v}_{\tt B}\cdot\nabla U_{\tt B})\Big),$ (69)
where $U_{\tt AB}=U_{\tt B}-U_{\tt A}$, $\mathbf{R}_{\tt AB}=\mathbf{r}_{\tt B}-\mathbf{r}_{\tt A}$, $R_{\tt AB}=|\mathbf{R}_{\tt AB}|$, $\mathbf{N}_{\tt AB}=\mathbf{R}_{\tt AB}/R_{\tt AB}$, $v_{\tt AB}=\mathbf{v}_{\tt A}-\mathbf{v}_{\tt B}$, and $b_{\tt B}=d(\mathbf{N}_{\tt AB}\cdot\mathbf{a}_{\tt B})/dt$. These expressions are consistent with our light-time Eq. (64) and close the model for range-rate and frequency transfer at the level needed for $10^{-16}$ work.
@@ III.2.2 Equations of motion in the GCRS
The trajectory of an Earth satellite in GCRS is governed by the following equation of motion [30, 1, 39, 15]:
$\ddot{\boldsymbol{\mathrm{r}}}(t)$ $=$ $\ddot{\boldsymbol{\mathrm{r}}}_{\tt E}+\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E}J_{2}}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt tidal}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt GR}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt SRP}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}elast}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}rad}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt custom},$ (75)
@@
$\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E},J_{2}}=$ $\frac{3J_{2,\tt E}\,\mu_{\tt E}R_{\tt E}^{2}}{2\,c^{2}\,r^{4}}\Big(\big[5\,\hat{\mathbf{r}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})^{2}-2\,\hat{\mathbf{S}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})-\hat{\mathbf{r}}\big]\big[v^{2}-\frac{4\mu_{\oplus}}{r}\big]$-
@@
In the worst-case orientation and using $J_{2,\oplus}=1.08263\times 10^{-3}$, $R_{\oplus}=6378~\mathrm{km}$, $\mu_{\tt E}=3.986\times 10^{14}~\mathrm{m^{3}s^{-2}}$, this acceleration was evaluated to contribute: LEO: $\sim 7\times 10^{-11}$ m s-2, MEO: $\sim 10\times 10^{-13}$ m s-2, cislunar: $\sim 5\times 10^{-16}$ m s-2. Thus, for LEO its omission biases precise orbit and gravity solutions; for MEO it is small but systematic; for cislunar it is negligible for dynamics but enter clock/frequency transfers only via $U$ coupling already modeled.
@@
Thus, relativistic contributions are important and must be kept in the model.
@@ III.2.3 Transforming equations of motion from GCRS to BCRS
Accurately calculating third-body acceleration is challenging due to the limitations of the GCRS in modeling interactions with celestial bodies. The procedure, as outlined in [1], Sec. 4.5.1, involves the following steps:
  • Convert the integration time from TT to TDB along the spacecraft’s worldline. As the difference between TT and TDB is periodic and typically less than 2 ms, this correction is generally negligible.
@@ III.3 Practical implementation and validation
In this Section, we outline the validation tests performed to verify the proper implementation of the EIH equations within the interplanetary orbit determination system used by GNSS spacecraft. The testing involved comparing the accelerations computed in the BCRS with those calculated in the GCRS. To enable this comparison, the accelerations derived in the GCRS were transformed into the BCRS using a series of relativistic transformations described in Sec. II.4. This approach ensures that the system accurately incorporates the relativistic effects necessary for precise navigation and tracking of GNSS spacecraft.
@@
Span 24 h; start epoch UTC (specified in run log)
@@
  • Imported and used the JPL planetary ephemeris DE440 for high-precision celestial mechanics.
@@
This testing procedure ensures that the implemented models and transformations meet the stringent requirements necessary for sub-millimeter accuracy in GNSS spacecraft navigation.
The numerical analysis conducted between the GCRS and the BCRS demonstrates that orbit states and observables can be precisely aligned, provided that the appropriate space-time transformations and relativistic corrections are applied. In the barycentric system, it is crucial not to overlook the relativistic corrections to acceleration. While in the Earth-centric system, these relativistic corrections to acceleration are less critical, the necessity for space-time transformations for conversion to and from the BCRS remains significant. In our testing scenario, after implementing the relativistic corrections and space-time transformations between the two systems, the discrepancy was observed to be under a millimeter in all three directions, see Fig. 1 within one day.
@@ IV Conclusions and next steps
We conducted a comprehensive analysis of general relativistic effects on the dynamics of near-Earth satellites within the GCRS and BCRS frameworks. Our results demonstrate that, when full relativistic corrections are applied, both reference systems yield consistent and equivalent results, enabling high-precision satellite navigation. These corrections are essential for achieving sub-centimeter accuracy in orbit determination and picosecond-level precision in time synchronization, which are especially critical for long-duration and deep-space missions, e.g., [47]. Our analysis supports the use of GNSS signals for precise navigation and timing within the Earth-Moon system, laying the foundation for future cislunar operations where cm-level accuracy in spatial positioning is required over large spatial regions.
Upcoming cislunar missions will necessitate a similar relativistic approach to ensure accurate time and position transformations between the BCRS, Earth, and the Moon. This calls for the development of a Lunicentric Coordinate Reference System (LCRS), analogous to the GCRS (see discussion in Appendix A and [9, 19].) A key step will be establishing a TDB-compatible time scale for the Moon, such as TCL (Temps Coordonné Lunaire), to synchronize lunar timekeeping with the BCRS. Additionally, the spatial scale in lunar coordinates must incorporate relativistic effects, such as Lorentz contraction, to meet the sub-millimeter accuracy demands of future lunar missions.
@@
By executing these demonstrations—first in MEO, then on the lunar surface, and finally via spacecraft in cislunar space—we will close the loop on both navigation and timing, fully leveraging GNSS signals across the Earth–Moon system. This LCRS deployment roadmap paves the way for cm-accurate cislunar navigation and ps-synchronized operations, enabling reliable guidance for Artemis, commercial landers, and future deep-space exploration.
@@ A.1 Introducing the LCRS
To define the Lunicentric Coordinate Reference System (LCRS), we will adopt a similar approach to that used for the GCRS. Following the structure of the GCRS metric [19], the LCRS with coordinates ($\cal{T}$ = TCL, $\boldsymbol{\mathrm{\cal{X}}}$) is described by a metric of similar form. The components of this metric are tailored to account for the Moon’s gravitational field, tidal effects from nearby massive bodies, and the relativistic corrections required for high-precision trajectory modeling can be expressed in the form (see details in [9]):
@@ A.5 Expressions to transform position, velocity and acceleration
$(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt M})_{\tt TDB}$ $=$ $\dot{\boldsymbol{\mathrm{\cal X}}}_{\tt TL}-c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+(1+\gamma)U_{\tt M,\rm ext}+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{v}}_{\tt M}\Big\}_{\tt TL}+{\cal O}\Big(7.82\times 10^{-9}~{\rm m/s}\Big),$~~~~~~ (109)
@@
$+$ $\Big((\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{a}}_{\tt M})+(1+2\gamma)\dot{U}_{\tt M,\rm ext}+4\big(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}})+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt M}+$
$+$ $\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\boldsymbol{\mathrm{\cal X}})+(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-(\dot{\boldsymbol{\mathrm{\cal X}}}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-({\boldsymbol{\mathrm{\cal X}}}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt M}+(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\boldsymbol{\mathrm{\cal X}}\Big\}_{\tt TL}+{\cal O}\Big(3.99\times 10^{-15}~{\rm m/s}^{2}\Big),\$
@@
Appendix B Reference transform test vectors (for code validation)
Table 5 provides a fixed self-consistent test set for code validation. States are referenced to DE440 (TDB argument) and ITRF2020-to-GCRS realized by IERS (2010), [5]. Use double precision (or higher) and report agreement to the thresholds quoted below each equation.
[…]
Item Input Output Notes
[…]

### v2

@@ II.6.2 Inverse transformations: proper time $\rightarrow$ TDB
III BCRS implementation and frame-closure validation for GNSS dynamics
[…]
For completeness, we summarize the equations of motion used in each frame and highlight the treatment of the dominant Newtonian, post-Newtonian, and geopotential terms needed to reproduce the closure experiment.
@@ III.1.1 BCRS metric in the mass monopole approximation
where the indices $j$ and $k$ refer to the $N$ bodies, with $k$ including body $i$, whose motion is being investigated. The gravitational constant for body $j$, denoted as $\mu_{j}$, is defined as $\mu_{j}=Gm_{j}$, where $G$ is the universal Newtonian gravitational constant, and $m_{j}$ is the isolated rest mass of body $j$. Additionally, the vector $\boldsymbol{\mathrm{r}}_{i}$ represents the barycentric radius-vector of this body. The vector $\boldsymbol{\mathrm{r}}_{ij}=\boldsymbol{\mathrm{r}}_{j}-\boldsymbol{\mathrm{r}}_{i}$ points from body $i$ to body $j$, $r_{ij}=|\boldsymbol{\mathrm{r}}_{j}-\boldsymbol{\mathrm{r}}_{i}|$, and the vector $\boldsymbol{\mathrm{n}}_{ij}=\boldsymbol{\mathrm{r}}_{ij}/r_{ij}$ is the unit vector along this direction.
@@ III.1.3 Light time equation
$t_{2}-t_{1}=\frac{r_{12}}{c}+(1+\gamma)\sum_{\tt B}\frac{GM_{\tt B}}{c^{3}}\ln\left[\frac{r_{1}^{\tt B}+r_{2}^{\tt B}+r_{12}}{r_{1}^{\tt B}+r_{2}^{\tt B}-r_{12}}\right]+\mathcal{O}(c^{-5}),$ (64)
where $t_{1}$ and $t_{2}$ refer to the signal’s emission and reception coordinate times, respectively, $\boldsymbol{\mathrm{r}}_{1}\equiv\boldsymbol{\mathrm{r}}(t_{1})$ and $\boldsymbol{\mathrm{r}}_{2}\equiv\boldsymbol{\mathrm{r}}(t_{2})$ are the barycentric position vectors of emitter and receiver, and $r_{12}\equiv|\boldsymbol{\mathrm{r}}_{2}-\boldsymbol{\mathrm{r}}_{1}|$. For each gravitating body $\tt B$} in the sum, $r_{1}^{\tt B}\equiv|\boldsymbol{\mathrm{r}}_{1}-\boldsymbol{\mathrm{x}}_{\tt B}(t_{1})|$ and $r_{2}^{\tt B}\equiv|\boldsymbol{\mathrm{r}}_{2}-\boldsymbol{\mathrm{x}}_{\tt B}(t_{2})|$ are the emitter/receiver distances to $\tt B$}. The Shapiro term is proportional to $(1+\gamma)$, with $\gamma=1$ in general relativity.
@@
III.1.4 Frequency transfer observable at $\mathcal{O}(c^{-3})$
@@
where the ratio $q_{A}/q_{B}$ follows from differentiating the light-time equation (64) with respect to emission time, retaining terms through $\mathcal{O}(c^{-3})$ [41, 10].
For GNSS applications the relevant observable is the one-way downlink. An explicit post-Newtonian expansion of the propagation factor $q_{A}/q_{B}$ through $\mathcal{O}(c^{-3})$, consistent with Eq. (64), is given in Ref. [41] (see also Refs. [1, 39]). In our implementation we compute $(u^{0})_{A}$ and $(u^{0})_{B}$ from the adopted metric and obtain $q_{A}/q_{B}$ from the derivative $\mathrm{d}(t_{2}-t_{1})/\mathrm{d}t_{1}$ evaluated along the modeled light-time solution. These expressions are consistent with our light-time (64) and close the model for range-rate and frequency transfer at the level needed for $10^{-16}$ work.
@@ III.2.2 Equations of motion in the GCRS
The trajectory of an Earth satellite in GCRS is governed by the following equation of motion [32, 1, 40, 43, 13]:
$\ddot{\boldsymbol{\mathrm{r}}}$ $=$ $\ddot{\boldsymbol{\mathrm{r}}}_{\tt E}+\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E}J_{2}}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt tidal}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt GR}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt SRP}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}elast}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}rad}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt custom},$ (73)
@@
$\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E},J_{2}}=$ $\frac{3J_{2,\tt E}\,\mu_{\tt E}R_{\tt E}^{2}}{2\,c^{2}\,r^{4}}\Big(\big[5\,\hat{\mathbf{r}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})^{2}-2\,\hat{\mathbf{S}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})-\hat{\mathbf{r}}\big]\big[v^{2}-\frac{4\mu_{\tt E}}{r}\big]$-
@@
In the worst-case orientation and using $J_{2,{\tt E}}=1.08263\times 10^{-3}$, $R_{\tt E}=6378~\mathrm{km}$, $\mu_{\tt E}=3.986\times 10^{14}~\mathrm{m^{3}s^{-2}}$, this acceleration was evaluated to contribute: LEO: $\sim 7\times 10^{-11}$ m s-2, MEO: $\sim 10\times 10^{-13}$ m s-2, cislunar: $\sim 5\times 10^{-16}$ m s-2. Thus, for LEO its omission biases precise orbit and gravity solutions, while for MEO it is small but systematic. A useful scaling is $\ddot{\boldsymbol{\mathrm{r}}}_{\tt EJ2\,rel}\propto r^{-4}$, so the effect rapidly decreases for higher Earth orbits. As a rough displacement scale over a one-day arc, $\tfrac{1}{2}|\ddot{\boldsymbol{\mathrm{r}}}|\,\Delta t^{2}$ with $\Delta t=86400$ s gives $\sim 0.3$ m (LEO, using $7\times 10^{-11}\,{\rm m/s^{2}}$) and $\sim 4$ mm (GPS-like MEO, using $1\times 10^{-12}\,{\rm m/s^{2}}$); at lunar distance it falls below the micron level. Accordingly, the 1PN $J_{2}$ term is mandatory for LEO POD, can matter for mm-level MEO processing, and is negligible for lunar-distance dynamics.
@@
For GNSS MEO orbit determination, planetary tidal terms beyond Sun/Moon are typically negligible at the cm level over daily arcs and are largely absorbed by empirical parameters; Eq. (77) quantifies that even worst-case Jupiter/Saturn tidal accelerations remain $\lesssim 10^{-11}$ m/s2.
@@
The de Sitter term written here is the dominant solar contribution for Earth satellites. Quantitatively, the dominant contribution to the geodetic (de Sitter) term in Eq. (78) is solar. For a near-Earth satellite, a simple magnitude estimate gives
$\frac{a^{\rm Moon}_{\rm dS}}{a^{\rm Sun}_{\rm dS}}\sim\frac{GM_{\tt M}\,v_{\tt EM}/r_{\tt EM}^{2}}{GM_{\tt S}\,v_{\tt ES}/r_{\tt ES}^{2}}\approx 2\times 10^{-4},$
where $(\boldsymbol{\mathrm{r}}_{\tt ES},\boldsymbol{\mathrm{v}}_{\tt ES})$ and $(\boldsymbol{\mathrm{r}}_{\tt EM},\boldsymbol{\mathrm{v}}_{\tt EM})$ denote the Earth–Sun and Earth–Moon relative state vectors. Thus, if the solar de Sitter acceleration is $\sim 10^{-11}$ m/s2 (81), the corresponding Moon term is $\lesssim 10^{-15}$ m/s2, and other planets are smaller still.
Thus, relativistic contributions presented above are important and must be kept in the model.
@@ III.2.3 Transforming equations of motion from GCRS to BCRS
Accurate evaluation of third-body accelerations in a TT-compatible GCRS propagation requires consistent use of barycentric ephemerides (typically delivered in TDB-compatible form) and the corresponding $T\leftrightarrowTDB$ time mapping. The procedure, as outlined in [1], Sec. 4.5.1, involves the following steps:
[…]
@@ III.3 Practical implementation and validation
In this Section, we outline the validation tests performed to verify the proper implementation of the EIH equations within the interplanetary orbit determination system used by GNSS spacecraft. The testing involved comparing the accelerations computed in the BCRS with those calculated in the GCRS. To enable this comparison, the accelerations derived in the GCRS were transformed into the BCRS using a series of relativistic transformations described in Sec. II.4. This approach ensures that the system accurately incorporates the relativistic effects necessary for precise navigation and tracking of GNSS spacecraft. Table 5 shows configuration used for the closure test.
We define the frame-closure residuals as
$\Delta\boldsymbol{\mathrm{r}}(T)\equiv\boldsymbol{\mathrm{X}}_{{\tt BCRS}\rightarrow{\tt GCRS}}(T)-\boldsymbol{\mathrm{X}}_{\tt GCRS}(T),\qquad\Delta\dot{\boldsymbol{\mathrm{r}}}(T)\equiv\dot{\boldsymbol{\mathrm{X}}}_{{\tt BCRS}\rightarrow{\tt GCRS}}(T)-\dot{\boldsymbol{\mathrm{X}}}_{\tt GCRS}(T),$ (82)
where $\boldsymbol{\mathrm{X}}_{\tt GCRS}(T)$ is obtained by integrating the GCRS equations of motion, and $\boldsymbol{\mathrm{X}}_{{\tt BCRS}\rightarrow{\tt GCRS}}(T)$ is obtained by transforming the initial state into the BCRS, integrating the corresponding barycentric equations of motion, and transforming the state back into the GCRS at each epoch. Figure 1 shows the components of $\Delta\boldsymbol{\mathrm{r}}(t)$ in the local radial/along-track/cross-track frame.
@@
Span 24 h; start epoch UTC: 2025-01-01 00:00:00
@@
  • Imported and used the JPL planetary ephemeris DE440 [4] for high-precision celestial mechanics.
@@
This testing procedure ensures that the implemented models and transformations meet the requirements for millimeter-level consistency in GNSS-enabled spacecraft navigation.
The numerical analysis conducted between the GCRS and the BCRS demonstrates that orbit states and observables can be precisely aligned, provided that the appropriate space-time transformations and relativistic corrections are applied. In the barycentric system, it is crucial not to overlook the relativistic corrections to acceleration. While in the Earth-centric system, these relativistic corrections to acceleration are less critical, the necessity for space-time transformations for conversion to and from the BCRS remains significant.
The frame-closure residuals defined in Eq. (82) compare the BCRS state obtained by direct integration in the BCRS to the BCRS state obtained by integrating in the GCRS and transforming the resulting state at the same epoch. For the configuration in Table 5, the residuals remain in the millimeter regime over the 24 h test interval (Fig. 1), with the largest excursion occurring in the along-track component (peak-to-peak $\sim$7 mm).
@@ IV Conclusions and next steps
[…]
Several of the necessary transformations, particularly between the GCRS and BCRS, have already been implemented in the JPL programs used to generate planetary and lunar ephemerides [38, 25], including the use of astronomical constants compiled before the GRAIL mission [55]. These systems utilize astronomical constants established prior to the GRAIL mission [55] but will require updates to account for increased accuracy demands. These models were also implemented in the high-precision tools like JPL’s GipsyX software, which provides centimeter-level accuracy in real-time satellite navigation and precise orbit determination. Future systems will need to accommodate relativistic time delays on the order of tens of picoseconds and positional discrepancies on the order of millimeters to ensure accurate navigation and timing, especially in lunar and deep-space environments.
@@
By executing such demonstrations—first in MEO and then in representative Earth–Moon geometries—one can validate the end-to-end GCRS/BCRS/LCRS reference-system infrastructure (state transformations, time scales, and 11PN light-time) in a mission-relevant configuration. This infrastructure is a prerequisite for cm-class cislunar navigation and ps-level time-transfer; achieved user-level performance will additionally depend on signal availability, measurement quality, and the estimation strategy.
@@ A.1 Introducing the LCRS
The International Astronomical Union has adopted a standard Lunar Celestial Reference System (LCRS) and Lunar Coordinate Time (TCL) [56]. In this work, we follow that framework and the detailed development in [21] to provide implementation-ready coordinate and time transformations between the BCRS and the LCRS. The LCRS is realized as a kinematically non-rotating reference system centered at the Moon and aligned with the ICRS axes, analogous to the relation between the BCRS and the GCRS. The realization of lunar body-fixed orientation and cartographic elements follows the IAU Working Group recommendations; see, e.g., [57]. These conventions enter the LCRS through the adopted lunar rotation model and the definition of surface site coordinates.
Following the structure of the GCRS metric [21], the LCRS with coordinates ($\cal{T}$ = TCL, $\boldsymbol{\mathrm{\cal{X}}}$) is described by a metric of similar form. The components of this metric are tailored to account for the Moon’s gravitational field, tidal effects from nearby massive bodies, and the relativistic corrections required for high-precision trajectory modeling can be expressed in the form (see details in [8]):
@@ A.5 Expressions to transform position, velocity and acceleration
$(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt M})_{\tt TDB}$ $=$ $\dot{\boldsymbol{\mathrm{\cal X}}}_{\tt TL}-c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+(1+\gamma)U_{\tt M,\rm ext}+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})\boldsymbol{\mathrm{v}}_{\tt M}\Big\}_{\tt TL}+{\cal O}\Big(7.82\times 10^{-9}~{\rm m/s}\Big),$~~~~~~ (108)
@@
$+$ $\Big((\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{a}}_{\tt M})+(1+2\gamma)\dot{U}_{\tt M,\rm ext}+4\big(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt M}+$
$+$ $\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\boldsymbol{\mathrm{\cal X}})+(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-(\dot{\boldsymbol{\mathrm{\cal X}}}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-({\boldsymbol{\mathrm{\cal X}}}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt M}+(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\boldsymbol{\mathrm{\cal X}}\Big\}_{\tt TL}+{\cal O}\Big(3.99\times 10^{-15}~{\rm m/s}^{2}\Big),$